### v2

@@ 1 Introduction
Figure 1: A qualitative comparison between the baseline and the improved segmentation with the tricks we propose (Ours). The first row shows images sampled from DUTS-TE [33], the second row shows the baseline segmentation predictions, the third row shows the segmentation results with our tricks, and the last row shows the ground truth segmentation masks. In (a) we show 4 image samples where we achieve a significant improvement over the baseline; in (b) we show two image samples where we achieve the same results as in the baseline; in (c) we show two image samples where the tricks make the segmentation masks worse (relative to the ground truth mask). We point out that some of the incorrect masks may also be due to the inherent ambiguity of the unsupervised segmentation task.
@@
Despite encouraging results, these methods seem to be either limited in resolution [23] or to rely on difficult training schemes [3]. Since many of the existing methods [21, 34, 23] rely on DINO features [5], which reduce the initial image resolution by a factor of $8$ or $16$, we introduce two rather general “tricks” to enhance the resolution of predicted segmentation maps and demonstrate their capabilities on the recent state of the art method Sempart for unsupervised segmentation learning [23]. [^1] https://github.com/alpErenSari/segmentation-tricks
A relatively naive attempt to enhance the resolution of the predicted segmentation masks is to simply compute higher-resolution DINO features and then plug them in the affinity matrix used in normalized cuts [23]. However, this immediately comes at a substantial computational cost (the complexity of normalized cuts scales cubically with the size of the affinity matrix) and thus quickly becomes numerically unfeasible.
As mentioned earlier on, we introduce two practical tricks. Our first trick is to use post-processing to enhance the resolution of the output masks. We show that well-established image processing methods such as guided filtering [15] can easily improve the accuracy of the predicted segmentation masks when the input image luminance is used as the filter guidance. Our second trick is a general criterion to enhance the resolution of dense predictions. We introduce a multi-scale consistency criterion on the predicted segmentation mask. We do so by using a teacher-student training scheme, where the teacher network takes a zoomed-in region of an image as input and the student network takes the whole image as input. Then, we zoom in on the same region of the predicted segmentation mask from the student network’s prediction and match it to the prediction from the teacher. Because we use the output of the teacher network as a target, we only backpropagate through the student network. A qualitative comparison between the baseline and our method can be seen in Fig. 1.
@@ 2 Related Work
Segmentation learning is a task with a long history in computer vision. Some of the classic methods (e.g., normalized cuts) find their way back into the training of deep neural networks [23]. Therefore, we revise some of these basic techniques, give a brief overview of supervised methods, and finally delve deeper into the unsupervised learning methods that relate the most to our techniques.
@@
Supervised Image Segmentation. Supervised segmentation methods have been developed on a variety of deep architectures [24, 6, 8, 22] and have recently demonstrated astonishing performance [17, 16]. However, because supervision requires lots of human labor to obtain accurate segmentation masks, these methods face remarkable challenges when they have to be scaled to large datasets.
Self-supervised representations (SSL). The idea of training deep networks without any supervision to obtain effective image representations gained popularity with the success of [7, 14, 4, 5, 13]. In particular, DINO [5] features based on ViT [9] demonstrate capabilities beyond image classification. As discussed in the next paragraph, recent methods for unsupervised segmentation learning find substantial gains by incorporating these SSL features.
@@
In this work, we extend the capabilities of Sempart [23], leveraging its recent advancements as the state-of-the-art (SotA) method in unsupervised segmentation learning. Despite its strengths, Sempart encounters limitations in resolution, primarily stemming from its reliance on low-resolution self-supervised learning (SSL) features for feedback. Our proposed tricks address this challenge in a practical and effective manner. These tricks are rather general and can be applied to other methods where an increase in resolution and accuracy is beneficial.
3 Background
@@
where $\mathbf{1}$ is a vector of ones. Eq. 5 is optimized only with respect to the low resolution mask $S_{\text{coarse}}$. High-resolution mask $S_{\text{fine}}$ is guided through a consistency loss
@@
4 Two Tricks
Recent unsupervised segmentation methods face diverse challenges, such as low-resolution feedback [21, 34, 23], instability during training due to the incorporation of adversarial losses [20, 2, 3], and misalignment between the predicted mask edges with the actual object edges. Many methods attempt to address the latter by employing bilateral filtering [1] as a post-processing step, though this approach has shown performance degradation in certain instances, as in [3, 23]. To tackle this issue, we propose to use guided filtering [15] as an alternative post-processing step, which has not been explored extensively in recent literature. Our experiments in Sec. 5 demonstrate that guided filtering significantly improves segmentation performance across various unsupervised methods.
The second trick we propose aims to mitigate the challenges associated with low-resolution feedback in methods incorporating SSL features. We introduce an equivariance loss based on random cropping of the input image and mask prediction in a teacher-student training scheme. The primary motivation behind this approach is to derive a finely detailed segmentation mask from a cropped and zoomed-in image, encouraging the student network to generate a matching prediction for the original image. We use the same networks as the student and the teacher. This process is illustrated in Fig. 3.
Guided Filtering. The guided filter [15] is a member of the edge-preserving smoothing image filter family. Although bilateral filtering [1] is commonly used in segmentation applications, we consider guided filtering because of two essential advantages compared to bilateral filtering: a lower computational complexity and no gradient reversal effects.
The guided filter is a locally linear filter with the hypothesis that the output image is an affine transformation of the input and this affine transformation is constant within a local window. Hence, the input and output pixel values can be expressed as $q_{i}=a_{k}I_{i}+b_{k}=p_{i}-n_{i},\forall i\in\omega_{k}$, where the $q_{i}$ is the output image, $p_{i}$ is the input image, $n_{i}$ is the estimated noise, $I_{i}$ is the guidance image, $a_{k}$ and $b_{k}$ are the affine transformation parameters, and $\omega_{k}$ is the local window. To estimate the parameters $a_{k}$ and $b_{k}$ the loss $E(a_{k},b_{k})=\sum\limits_{i\in\omega_{k}}((a_{k}I_{i}+b_{k}-p_{i})^{2}+\epsilon{a_{k}}^{2})$ is minimized for all local windows $\omega_{k}$. After taking the derivative of the loss, and equating it to 0, we get $a_{k}=(\frac{1}{|\omega|}\sum_{i\in\omega_{k}}I_{i}p_{i}-\mu_{k}\bar{p}_{k})/(\sigma_{k}^{2}+\epsilon)$ and $b_{k}=\bar{p}_{k}-a_{k}\mu_{k}$. We will denote the guided filtering process as $\text{GF}(\cdot,\cdot)$ in the rest of the paper. For instance, $q=\text{GF}(p,I)$ means that the image $p$ is filtered with the image $I$’s guidance, and we obtain $q$ as the filtered image.
Figure 2: Some sampled images from DUTS-TE [33] containing small or intricate objects, which are challenging for segmentation tasks.
We propose to set $p$ as the predicted segmentation mask and $I$ as the luminance of the input image. In other words, we apply guided filtering on the predicted mask with the guidance of the grayscale version of the input image. Since guided filtering is essentially an affine transformation, the gradients of the filtered image and the guidance image will be proportional, which can be regarded as the alignment of the input image edges according to the guidance from the image edges. Multi-Scale Consistency. One challenge associated with incorporating a segmentation head over a ViT [9] feature map (as with DINO) is the loss of details, particularly for small objects or structures within the image. Sempart [23] addresses this issue by feeding the input image into the segmentation head and by applying a total variation loss on the resulting high-resolution mask. Despite these efforts, the model suffers from inaccurate segmentations of small objects or of objects with thin structures, as illustrated in Fig. 2. We aim to tackle this problem by introducing an equivariance loss based on random cropping. Whereas Leopart [39] previously proposed a similar concept for training self-supervised image representations, we extend this idea by directly applying the equivariance loss to the segmentation prediction, as depicted in Fig. 3. The proposed equivariance loss is computed through the following steps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Apply stop gradient afterwards to prevent a collapse of the predicted masks, which we denote as $\text{SG}[\text{GF}(\hat{s}_{c})]=s_{target}$, where $s_{target}$ is the final target for our loss, and $\text{SG}[\cdot]$ is the stop gradient operation.
  5. 5.
In our experiments, we added $\mathcal{L}_{eq}$ multiplied with the parameter $\lambda_{eq}$ to the overall loss (10). We also used the cropped image to train the network.
5 Experimental Results
Table 1: The unsupervised saliency segmentation results on DUT-OMRON [36], DUTS-TE [33], and ECSSD [27]. The best results are indicated in bold. ${}^{\ast}$ indicates our implementation of Sempart [23] since the code is currently unavailable. “Ours” indicates the application of the proposed trick over the Sempart baseline. We demonstrate improvements over the baseline implementation we have.
Model DUT-OMRON [36] DUTS-TE [33] ECSSD [27]
Acc IoU $maxF_{\beta}$ Acc IoU $maxF_{\beta}$ Acc IoU $maxF_{\beta}$
LOST [29] .797 .410 .473 .871 .518 .611 .895 .654 .758
TokenCut [34] .880 .533 .600 .903 .576 .672 .918 .712 .803
FOUND [30] - single .920 .586 .683 .939 .637 .733 .912 .793 .946
MOVE [3] .923 .615 .712 .950 .713 .815 .954 .830 .916
Sempart [23] .932 .668 .764 .959 .749 .867 .964 .855 .947
${}^{\ast}$908 .664 .758 .949 .761 .856 .950 .853 .925
Ours .920 .684 .765 .956 .787 .869 .952 .862 .931
MOVE [3] + BF .931 .636 .734 .951 .687 .821 .953 .801 .916
Sempart + BF .942 .698 .799 .958 .749 .879 .963 .850 .944
${}^{\ast}$ + BF .914 .689 .770 .950 .765 .857 .950 .849 .925
Ours + BF .920 .675 .763 .950 .742 .855 .945 .830 .920
SelfMask w/ MOVE .933 .666 .756 .954 .728 .829 .956 .835 .921
SelfMask w/ Sempart .942 .698 .799 .958 .749 .879 .963 .850 .944
SelfMask w/ ${}^{\ast}$923 .674 .768 .953 .743 .866 .962 .854 .942
SelfMask w/ Ours .935 .685 .790 .956 .747 .881 .962 .852 .948
5.1 Implementation
Since the code of Sempart [23] has not been released yet, we implemented it by following the implementation details in the paper with minor changes. The Sempart [23] results obtained with our implementation are denoted as “${}^{\ast}$ [^2] The implementation can be found at https://github.com/alpErenSari/segmentation-tricks
Sempart. We utilize the ViT-s/8 version of the official DINO [5], set input image size to $320\times 320$, extract the keys from the last layer with size $384\times 40\times 40$, and use them to perform segmentation and compute the normalized cut loss. AdamW [19] is utilized for optimization with a learning rate of $1\times{10}^{-4}$, $\beta=(0.9,0.99)$, and weight_decay of 0. Everything was implemented in Pytorch and trained for 30 epochs on a single NVIDIA GTX 3090 GPU with a batch size of 8. We set $\lambda_{SR}=1$, $\lambda_{\text{GTV-coarse}}=1$, $\lambda_{\text{GTV-fine}}=1$, $\lambda_{\text{Ncut-Crop}}=0.01$, and $\lambda_{eq}=0.01$ where $\lambda_{\text{Ncut-Crop}}$ is the normalized cut loss calculated on the random cropped and upsampled image region. For graph total variation losses, the average of each pixel’s loss was used instead of the summation, since we observed that this improves stability.
Data augmentation. The shorter dimension (width or height) of the input images are first resized to 355, and a random $320\times 320$ region is cropped. We also used slight color jittering and random grayscale with a probability of 0.2.
Graph affinity. Contrary to Sempart [23], we use keys from the last layer of ViT-S/8 DINO, and set $\tau=0.4$ and $\epsilon=1\times{10}^{-6}$. These changes were made since our implementation following the exact settings of Sempart failed to converge. We keep the hyper-parameters for $\mathcal{L}_{GTV-coarse}$ and $\mathcal{L}_{GTV-fine}$ same.
Foreground selection. During the evaluation, we calculate the average over all pixels in a batch, and we invert the masks if this mean is greater than 0.5. The assumption is that the foreground should cover a smaller area than the background on average.
Multi-scale consistency. We randomly crop the same $80\times 80$ fraction of the input image and the mask prediction. Afterward, these crops are upscaled to $320\times 320$. We also used the random cropped image to train the Sempart [23].
Guided filtering. To apply guided filtering on segmentation predictions, initially, we re-scale the pixel values of the guidance image into the range of 0 and 1, just like the mask prediction. Afterward, the pixel values of the guidance image are averaged to get a grayscale estimation of the image. This final one-channel image is utilized for the guided filtering.
5.2 Unsupervised Saliency Segmentation
Datasets. We used the DUTS-TR [33] dataset which contains 10,553 object-centric images for training in every experiment, as [28, 3, 23]. To evaluate the performance of unsupervised saliency segmentation, we used DUT-OMRON [36] dataset with 5,168 images, DUTS-TE [33] dataset with 5,019 images, and ECSSD [27] dataset with 1,000 images.
[…]
with $\beta=0.3$.
Results. In Tab. 1, we present a comprehensive comparison of our results with the recent state-of-the-art method Sempart [23]. Our implementation, denoted as “${}^{\ast}$”, undergoes evaluation across three distinct criteria: baseline results, post-processed baseline results employing bilateral filtering [1], and outcomes derived from SelfMask [28], trained with pseudo-masks generated by various methods. The comparison reveals the efficacy of our approach in advancing unsupervised saliency segmentation.
[…]
5.3 Single Object Detection
@@
Evaluation. First, the predicted mask is split into parts if it’s not fully connected. Then, we select the mask with the largest bounding box as our object prediction and calculate the IoU between our prediction and the ground-truth bounding box. Finally, the percentage of the IoUs higher than 0.5 is calculated, which is known as the Correct Localization (CorLoc) metric.
Results. In Tab. 2, we conduct a comprehensive comparison between our results and the recent state-of-the-art MOVE [3], as well as our baseline in object detection. Our tricks significantly enhance the baseline across all three datasets considered. This improvement is evident in detection accuracy, underscoring the effectiveness of our approach in advancing unsupervised object detection.
The observed disparities between the reported baseline and our implemented version prompts the assertion that our approach has the potential to establish a new state-of-the-art benchmark in both Pascal VOC07 [10] and COCO20K [18] datasets. This claim is substantiated by the consistent and noteworthy improvements achieved across various evaluation metrics.
Our ablation study in Sec. 5.4 provides additional insights, revealing that random cropping emerges as a key contributor to the overall improvement in object detection. Specifically, Fig. 4 visually depicts how random cropping facilitates the detection of small objects and, in some instances, objects with occlusion. This emphasizes the nuanced impact of each proposed technique, with random cropping playing a pivotal role in enhancing the model’s capability to handle challenging detection scenarios.
@@
Table 2: Comparison of single object detection on Pascal VOC07 [10], Pascal VOC12 [11], COCO20K [18] with CorLoc metric. The best results are indicated in bold. ${}^{\ast}$ indicates our implementation of Sempart [23]. ($\uparrow z$) indicate the improvements over the baseline.
Model       VOC07 [10]       VOC12 [11] COCO20K [18]
LOST [29] 61.9 64.0 50.7
Deep Spectral [21] 62.7 66.4 50.7
TokenCut [34] 68.8 72.1 58.8
MOVE [3] 76.0 78.8 66.6
Sempart [23] 75.1 76.8 66.4
${}^{\ast}$ 71.7 75.9 61.3
Ours          74.4 ($\uparrow 2.7$)          77.0 ($\uparrow 1.1$)          63.4 ($\uparrow 2.1$)
5.4 Ablations
@@
The tricks on saliency segmentation. In Tab. 3, we present the Intersection over Union (IoU) results obtained from the evaluation of unsupervised saliency segmentation on three datasets: DUT-OMRON [36], DUTS-TE [33], and ECSSD [27]. The original baseline is trained with DINO [5] ViT-S/8 as the backbone. Additionally, we provide results for baselines trained with DINO ViT-S/16 and IBOT [38] ViT-S/16 as the backbones. Our findings indicate that across different backbones, the proposed techniques significantly enhance segmentation performance. Notably, guided filtering emerges as a key contributor to the improved segmentation results.
@@
The tricks on single object detection. In Tab. 4, we demonstrate the enhancement achieved by our techniques in object detection based on the CorLoc metric. The evaluation is performed on datasets comprising the train and validation splits of Pascal VOC07 [10] and Pascal VOC12 [11], as well as the train split of COCO20K [18]. Notably, unlike the case of saliency segmentation, we observe that random cropping yields the most significant improvement in object detection performance.
[…]
Table 5: The ablations of guided filtering [15] and bilateral filtering [1] with TokenCut [34] and MOVE [3] on unsupervised saliency segmentation with same datasets and metrics in Tab. 1. The best results for each method are indicated with bold.
@@
5.5 Limitations and Potential Negative Impacts
Limitations. Our method exhibits challenges in scenarios where saliency within an image is unambiguous or when multiple objects are present in the scene. Some visual examples can be seen in Fig. 5, where our method tends to segment all objects in the scene, whereas human annotators are inclined to select smaller and easily movable objects. This discrepancy highlights a limitation in our model’s ability to prioritize and differentiate between objects based on their perceived saliency. Moreover, in instances where the salient object and the background share visual similarities our method encounters difficulty, making it challenging to distinguish between them accurately. This issue is particularly pronounced when the visual features of the salient and background objects closely resemble each other. The struggle to discern between these visually similar elements contributes to inaccuracies in segmentation, as depicted in various scenarios.
Potential negative impacts. Our method is unlikely to have negative impacts, as its primary function is segmenting objects in images, which can be accomplished by alternative methods or through manual annotation by humans.

### v3

@@ 1 Introduction
Despite encouraging results, these methods seem to be either limited in resolution [19] or to rely on difficult training schemes [3]. Since many of the existing methods [18, 28, 19] rely on DINO features [5], which reduce the initial image resolution by a factor of $8$ or $16$, we introduce two rather general “tricks” to enhance the resolution of predicted segmentation maps and demonstrate their capabilities on the recent state of the art method Sempart for unsupervised segmentation learning [19].  [^1] https://github.com/alpErenSari/segmentation-tricks Our first trick is to use post-processing to enhance the resolution of the output masks. We show that well-established image processing methods such as guided filtering [13] can easily improve the accuracy of the predicted segmentation masks when the input image luminance is used as the filter guidance.
Our second trick is a general criterion to enhance the resolution of dense predictions. We introduce a multi-scale consistency criterion on the predicted segmentation mask. We do so by using a teacher-student training scheme, where the teacher network takes a zoomed-in region of an image as input and the student network takes the whole image as input. Then, we zoom in on the same region of the predicted segmentation mask from the student network’s prediction and match it to the prediction from the teacher. Because we use the output of the teacher network as a target, we only backpropagate through the student network.
@@ 2 Related Work
Self-supervised representations (SSL). The idea of training deep networks without any supervision to obtain effective image representations gained popularity with the success of [6, 12, 4, 5, 11]. In particular, DINO [5] features based on ViT [7] demonstrate capabilities beyond image classification.
@@
3 Two Tricks
We introduce two tricks that can be incorporated into different unsupervised segmentation methods in a flexible manner: guided filtering[13] and random cropping-based equivariance loss.
Guided Filtering. First, the RGB image -whose saliency mask is predicted- is converted to grayscale. Afterward, the guided filter [13] is applied to the predicted segmentation mask using the grayscale image as guidance.
Multi-Scale Consistency. One challenge associated with incorporating a segmentation head over a ViT [7] feature map (as with DINO) is the loss of details, particularly for small objects or structures within the image. Sempart [19] addresses this issue by feeding the input image into the segmentation head and by applying a total variation loss on the resulting high-resolution mask. We aim to tackle this problem by introducing an equivariance loss based on random cropping depicted in Fig. 1.
The proposed equivariance loss is computed through the following steps:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Apply stop gradient afterward to prevent a collapse of the predicted masks, which we denote as $\text{SG}[\text{GF}(\hat{s}_{c})]=s_{target}$, where $s_{target}$ is the final target for our loss, and $\text{SG}[\cdot]$ is the stop gradient operation.
  5. 5.
In our experiments, we added $\mathcal{L}_{eq}$ multiplied with the parameter $\lambda_{eq}$ to the overall loss of Sempart [19]. We also used the cropped image to train the network.
4 Experimental Results
Table 1: The unsupervised saliency segmentation results on DUT-OMRON [30], DUTS-TE [27], and ECSSD [22]. The best results are indicated in bold. Sempart indicates our implementation of Sempart [19] since the code is currently unavailable. “Ours” indicates the application of the proposed trick over the Sempart baseline. We demonstrate improvements over the baseline implementation we have.
Model DUT-OMRON [30] DUTS-TE [27] ECSSD [22]
Acc IoU $maxF_{\beta}$ Acc IoU $maxF_{\beta}$ Acc IoU $maxF_{\beta}$
LOST [24] .797 .410 .473 .871 .518 .611 .895 .654 .758
TokenCut [28] .880 .533 .600 .903 .576 .672 .918 .712 .803
FOUND [25] - single .920 .586 .683 .939 .637 .733 .912 .793 .946
MOVE [3] .923 .615 .712 .950 .713 .815 .954 .830 .916
Sempart [19] .932 .668 .764 .959 .749 .867 .964 .855 .947
Sempart .908 .664 .758 .949 .761 .856 .950 .853 .925
Ours .920 .684 .765 .956 .787 .869 .952 .862 .931
MOVE [3] + BF .931 .636 .734 .951 .687 .821 .953 .801 .916
Sempart + BF .942 .698 .799 .958 .749 .879 .963 .850 .944
Sempart + BF .914 .689 .770 .950 .765 .857 .950 .849 .925
Ours + BF .920 .675 .763 .950 .742 .855 .945 .830 .920
SelfMask w/ MOVE .933 .666 .756 .954 .728 .829 .956 .835 .921
SelfMask w/ Sempart .942 .698 .799 .958 .749 .879 .963 .850 .944
SelfMask w/ Sempart .923 .674 .768 .953 .743 .866 .962 .854 .942
SelfMask w/ Ours .935 .685 .790 .956 .747 .881 .962 .852 .948
4.1 Unsupervised Saliency Segmentation
Datasets. We used the DUTS-TR [27] for training in every experiment, as [23, 3, 19]. To evaluate the saliency segmentation, we used DUT-OMRON [30], DUTS-TE [27], and ECSSD [22] datasets.
Evaluation. Three metrics are reported in our work as [23, 3, 19], which are pixel mask accuracy (Acc), intersection over union (IoU), and $\max\textit{F}_{\beta}$ [28].
Results. In Table 1, we present a comprehensive comparison of our results with the recent state-of-the-art method Sempart [19]. Our implementation, denoted as “Sempart”, undergoes evaluation across three distinct criteria: baseline results, post-processed baseline results employing bilateral filtering [1], and outcomes derived from SelfMask [23], trained with pseudo-masks generated by various methods. The comparison reveals the efficacy of our approach in advancing unsupervised saliency segmentation.
Our results show state-of-the-art Intersection over Union (IoU) performance across all three datasets. When considering the notable gaps between our implemented baseline and the reported baseline, particularly in terms of accuracy (Acc) and $\max\textit{F}_{\beta}$ metrics, it becomes evident that our approach has the potential to establish a new state-of-the-art benchmark with precisely the same baseline.
4.2 Single Object Detection
Evaluation. We follow the evaluation protocol of the previous methods [28, 3, 19] and report Correct Localization (CorLoc) metric.
Results. In Table 2, we conduct a comprehensive comparison between our results and the recent state-of-the-art MOVE [3], as well as our baseline in object detection. Our tricks significantly enhance the baseline across all three datasets considered. This improvement is evident in detection accuracy, underscoring the effectiveness of our approach in advancing unsupervised object detection.
Table 2: Comparison of single object detection on Pascal VOC07 [8], Pascal VOC12 [9], COCO20K [16] with CorLoc metric. The best results are indicated in bold. Sempart indicates our implementation of Sempart [19]. ($\uparrow z$) indicate the improvements over the baseline.
Model       VOC07 [8]       VOC12 [9] COCO20K [16]
LOST [24] 61.9 64.0 50.7
Deep Spectral [18] 62.7 66.4 50.7
TokenCut [28] 68.8 72.1 58.8
MOVE [3] 76.0 78.8 66.6
Sempart [19] 75.1 76.8 66.4
Sempart 71.7 75.9 61.3
Ours          74.4 ($\uparrow 2.7$)          77.0 ($\uparrow 1.1$)          63.4 ($\uparrow 2.1$)
5 Conclusion
We introduce two practical methods that can be flexibly incorporated into unsupervised image segmentation. We demonstrate their effectiveness with state-of-the-art results on unsupervised saliency segmentation tasks, combined with the recent state-of-the-art baseline. Moreover, in the appendix, we demonstrate that these two methods can be effective even if the baseline was trained with different backbones and combined with other recent unsupervised segmentation methods. We also include related background information, more qualitative results illustrating the gains of our methods, and an ablation study in unsupervised saliency segmentation and unsupervised object discovery in the appendix.
Appendix A Additional Background
@@
where $\mathbf{1}$ is a vector of ones. Eq. 5 is optimized only for the low-resolution mask $S_{\text{coarse}}$. High-resolution mask $S_{\text{fine}}$ is guided through a consistency loss
@@
Appendix B Additional Results
B.1 Unsupervised Saliency Segmentation
We provide some qualitative results in Fig. 2 for the cases where the segmentation is improved, unaffected, and degraded.
Figure 2: A qualitative comparison between the baseline and the improved segmentation with our proposed tricks (Ours). The first row shows images sampled from DUTS-TE [27], the second row shows the baseline segmentation predictions, the third row shows the segmentation results with our tricks, and the last row shows the ground truth segmentation masks. In (a) we show 4 image samples where we achieve a significant improvement over the baseline; in (b) we show two image samples where we achieve the same results as in the baseline; in (c) we show two image samples where the tricks make the segmentation masks worse (relative to the ground truth mask). We point out that some of the incorrect masks may also be due to the inherent ambiguity of the unsupervised segmentation task.
B.2 Single Object Detection
@@
We share the implementation details of the single object detection evaluation and some qualitative results here to illustrate how our tricks boost object detection. Specifically, Fig. 3 visually depicts how random cropping facilitates the detection of small objects and, in some instances, objects with occlusion. Evaluation Details. First, the predicted mask is split into parts if it’s not fully connected. Then, we select the mask with the largest bounding box as our object prediction and calculate the IoU between our prediction and the ground-truth bounding box. Finally, the percentage of the IoUs higher than 0.5 is calculated, known as the Correct Localization (CorLoc) metric.
@@
B.3 Ablations
@@
The tricks on saliency segmentation. In Table 3, we present the Intersection over Union (IoU) results obtained from the evaluation of unsupervised saliency segmentation on three datasets: DUT-OMRON [30], DUTS-TE [27], and ECSSD [22]. The original baseline is trained with DINO [5] ViT-S/8 as the backbone. Additionally, we provide results for baselines trained with DINO ViT-S/16 and IBOT [32] ViT-S/16 as the backbones. Our findings indicate that across different backbones, the proposed techniques significantly enhance segmentation performance. Notably, guided filtering emerges as a key contributor to the improved segmentation results.
@@
The tricks on single object detection. In Table 4, we demonstrate the enhancement achieved by our techniques in object detection based on the CorLoc metric. The evaluation is performed on datasets comprising the train and validation splits of Pascal VOC07 [8] and Pascal VOC12 [9], as well as the train split of COCO20K [16]. Notably, unlike the case of saliency segmentation, we observe that random cropping yields the most significant improvement in object detection performance.
Guided filtering with different methods. In Table 5, our findings indicate that the application of guided filtering significantly enhances the performance of recent saliency segmentation methods TokenCut [28] and MOVE [3]. We present results Acc, IoU, and the $max\textit{F}_{\beta}$ metric across evaluation datasets including DUT-OMRON [30], DUTS-TE [27], and ECSSD [22]. Comparative analysis with baseline methods, incorporating bilateral filtering [1], is also included. Notably, the results in Table 5 demonstrate consistent improvement in all metrics with guided filtering, which contrasts with the performance of bilateral filtering.
Table 5: The ablations of guided filtering [13] and bilateral filtering [1] with TokenCut [28] and MOVE [3] on unsupervised saliency segmentation with same datasets and metrics in Table 1. The best results for each method are indicated with bold.
@@
B.4 Limitations
Our method exhibits challenges in scenarios where saliency within an image is unambiguous or when multiple objects are present in the scene. Some visual examples can be seen in Fig. 4, where our method tends to segment all objects in the scene, whereas human annotators are inclined to select smaller and easily movable objects. Moreover, in instances where the salient object and the background share visual similarities our method encounters difficulty, making it challenging to distinguish between them accurately.